OpenStudy (anonymous):

Find the perimeter of the rectangle A rectangle with a length of 9x-5 and a width of 2x + 4 A. 11x + 9 B. 22x - 2 C. 18x - 20 D. 11x - 1

Directrix (directrix):

What did you get?

OpenStudy (anonymous):

C.

Directrix (directrix):

Rectangles have opposite sides congruent. The perimeter is 2(9x - 5) + 2 (2x + 4) = 18x - 10 + 4x + 8 = ? I don't think it is C. What do you think?

OpenStudy (anonymous):

B. ?

Directrix (directrix):

Just combine like terms here: 18x - 10 + 4x + 8 = 18x + 4x - 10 + 8 = ? Is that sum shaping up to look as if it is B?

Directrix (directrix):

18x + 4x - 10 + 8

Directrix (directrix):

And, the answer is ?

OpenStudy (anonymous):

i got b

Directrix (directrix):

>>>B. ? Yes or No.

Directrix (directrix):

That's correct.

OpenStudy (anonymous):

thank you

Directrix (directrix):

Glad to help.